OpenStudy (oppskbeetlejuice):

Find the percent of markup for $40 marked up to $59.75. Round to the nearest percent. A. 20% B. 27% C. 49% D. 33%

OpenStudy (tkhunny):

Insufficient information. Which kind of markup is being used? 59.75 - 40.00 = 19.75 19.75/40.00 = 49.38% 19.75/59.75 = 33.05%